OpenStudy (anonymous):

The sum of the ages of Andrea and her mother is 45. Fifteen years from now, the mother will be twice as old as Andrea. How old is each now?

OpenStudy (gorv):

35 and 10 initially and after 15 year 50 and 25 year old

OpenStudy (anonymous):

Howd you come up with that answer? @gorv

OpenStudy (gorv):

@jillang07 andrea + mother =45 after 15 years mother +15=2(andrea+15) mother=2*andrea+15 now putting mother in (andrea + mother=45) andrea+ 2*andrea+15=45 3*andrea=30 andrea=10 year aftr 15 yr andrea=10+15=25 mother=25*2=50